# Generalized Out-of-Distribution Detection and Beyond in Vision Language Model Era: A Survey

### v1

@@ 5.4.2 Similarity between Each Methodology
Each Problem Setting  The existing problem settings for CLIP-based AD and CLIP-based OOD detection are similar. Both primarily focus on zero-shot and few-shot settings and can be categorized into training-free, auxiliary training-based, and ID training-based methods. By examining each problem setting more closely, we can observe, for instance, that while open-vocabulary AD is predominant in CLIP-based AD, it has not been deeply explored in CLIP-based OOD detection. This provides valuable insights into future directions that are crucial for OOD detection.
@@
History of Approaches  The history of the progress of methods for CLIP-based AD and CLIP-based OOD detection is similar. For instance, both problems initially started with naive methods with manual OOD prompts (ZeroOE [38] for OOD detection, WinCLIP [27] for AD). To address the issues with these initial approaches, subsequent methods emerged that replaced OOD prompts with learnable parameters (LSN [44] and NegPrompt [47] for OOD detection, and AnomalyCLIP [29] for AD). Therefore, by carefully examining each other’s fields, there is potential for mutual enhancement and interaction in the future.
@@ 6 Evolution in LVLM Era
In this section, we introduce the early advances in OOD detection and AD in the Large Vision Language Models (LVLM) era. While previous sections focused on VLMs such as CLIP, this section shifts our focus to the more emerging topic of “Large” VLMs. Recent advancements in computer vision have led to the emergence of LVLMs such as GPT-4V [30] and LLaVA [31]. Although these fields are still in their early stages with limited papers, this survey provides a deep introduction to each problem in the hope that our detailed review can help foster further advancements in this area.
Figure 4: Overview of the evolution of each problem in the Large Vision Language Model (LVLM) era. (i) Sensory AD has become a highly active and noteworthy field in the LVLM era [142]. (ii) Notably, OOD detection is evolving into a new task called Unsolvable Problem Detection [143] in the LVLM era. Figure adapted from [142] and partially from [143].
@@ 6.1 Change of Each Problem
i. Sensory AD $\rightarrow$ Sensory AD  Sensory AD has continued to develop in the LVLM era [144, 142, 145]. The use of LVLMs has made AD applicable in many domains and modalities [142].
ii. OOD Detection $\rightarrow$ Unsolvable Problem Detection  In the LVLM era, OOD detection has evolved into a new task termed Unsolvable Problem Detection (UPD) [143]. UPD evaluates the LVLMs’ ability to recognize and abstain from answering unexpected or unsolvable input questions, effectively expanding the scope of OOD detection into the context of Visual Question Answering (VQA) tasks. This shift to the VQA task has significantly broadened the concept of OOD detection to a wider range of AI tasks involving LVLMs.
@@ 6.2.1 Summary of Problem
Background  Following the recent revolutionary development of LLMs [146, 147, 148, 149, 150, 151], LVLMs [152, 153, 154, 149, 155, 156, 157, 158, 159] have demonstrated remarkable capabilities in diverse applications [160, 161, 162, 163]. However, a significant concern has arisen regarding the reliability of these models, specifically their ability to generate accurate and trustworthy information. These models frequently produce incorrect or misleading information, a phenomenon referred to as “hallucination” [32]. Among the various hallucination issues [32], the challenge of identifying out-of-place questions is crucial for deploying LVLMs in safety-critical applications. This challenge extends the concept of OOD detection to the VQA tasks for LVLMs and represents a specific aspect of LVLMs’ trustworthiness.
@@
Definition  Unsolvable Problem Detection (UPD) is a task to measure the trustworthiness of LVLMs, which is designed to evaluate models’ capacity to withhold answers when faced with unsolvable problems. The UPD task can be categorized into three distinct problem types: Absent Answer Detection (AAD), Incompatible Answer Set Detection (IASD), and Incompatible Visual Question Detection (IVQD). The details of each setting are as follows:
  1. 1.
    Absent Answer Detection (AAD): AAD evaluates the model’s capacity to determine when the correct answer is absent from the provided options.
  2. 2.
    Incompatible Answer Set Detection (IASD): IASD assesses the model’s ability to discern answer choices that are completely irrelevant to the given question and image.
  3. 3.
    Incompatible Visual Question Detection (IVQD): IVQD evaluates the model’s capacity to discern whether a question and image are unrelated or mismatched.
@@
Benchmark  Miyai et al. [143] created MM-UPD Bench for the UPD challenge. MM-UPD encompasses MM-AAD, MM-IASD, and MM-IVQD benchmarks for each UPD problem. Each benchmarks are created on the top of MMBench (dev) [162], which is a systematically designed objective benchmark for evaluating various abilities of LVLMs. Following the definition of each ability in MMBench (e.g., “Coarse Perception: Image Scene” and “Logic Reasoning: Future Prediction”), MM-UPD evaluates the trustworthiness of LVLMs from various abilities.
@@
Although MM-UPD is the main benchmark, the adaptation cost of creating UPD problems is not high, making it highly applicable to other benchmarks. For instance, the recently proposed MuirBench [164], a benchmark for multi-image understanding, has incorporated the concept of UPD by adding unsolvable problems.
@@
Application  UPD has a wide range of applications, from everyday use of LVLMs to robot manipulation. Especially when incorporating LVLMs into safety-critical domains such as robot manipulation [165] and autonomous driving [166], there is a risk of significant problems if the LVLM fails to identify erroneous user questions and makes incorrect predictions. UPD serves as a task to ensure safety in such safety-critical scenarios.
Evaluation  UPD introduces new evaluation metrics that incorporate the concept of the evaluation protocols for OOD detection, taking into account the prediction distribution for both standard (ID) and UPD (OOD) samples. The rationale is that ideal LVLMs should not only give correct answers for the standard problems but need to withhold answering in the UPD scenario where the problem is unsolvable. To better reflect the ideal behavior of LVLMs, UPD measures several metrics: (i) Standard Accuracy: The accuracy on standard problems where the image, question, and answer sets are all aligned, and the ground-truth answer is always present within the provided options. (ii) UPD Accuracy: The accuracy of AAD/IASD/IVQD problems. (iii) Dual Accuracy: Accuracy of standard and UPD pairs. We count success only if the model is correct for both standard and UPD problems.
@@ 6.2.2 Findings
In the following, we briefly summarized the findings of the UPD challenge [143].
1. Most LVLMs Hardly Hesitate to Answer. Most LVLMs, especially open-source LVLMs, have significantly low UPD accuracies, which indicates the difficulty of the UPD challenge. For example, LLaVA-1.5 [31] and CogVLM [155], which are state-of-the-art LVLMs, completely fail to withhold answering. GPT-4V achieves higher performances than other LVLMs due to its safety training process [167]. However, there is still a performance gap from the upper bound scores.
@@ 6.3.1 Summary of Problem
Background  Anomaly detection is a crucial task in a variety of domains and data types. However, existing anomaly detection models are often designed for specific domains or modalities [142]. Also, current AD methods only provide an anomaly score for the test sample and require a manual threshold to distinguish between normal and anomalous instances for each sample [144]. To facilitate real-world applications, developing a system capable of expressing anomalies in natural language across various modalities and domains is crucial for ensuring accessibility to a wider range of users.
Definition  The definition of AD remains consistent with conventional and CLIP-based AD, aiming to identify samples that deviate from predefined normality. The key distinction lies in the output. While previous methods produced an anomaly score requiring a manual threshold, AD with LVLMs aims to recognize and describe anomalies using text, eliminating the need for manual thresholds and enhancing human interpretability.
Benchmark  Since the field of AD with LVLMs is in its infant stage, there are still no unified benchmarks. AnomalyGPT [144] focuses on industrial image anomaly detection/localization and uses the standard benchmarks MVTec-AD [52] and VisA [53]. More recently, Cao et al. [142] extend the domain and modality and demonstrate the applications in industrial image anomaly detection/localization (e.g., MVTec-AD [52]), point cloud anomaly detection (MVTec 3D [168]), medical image anomaly detection/localization (e.g., Chest X-ray [169], Head CT [170]), logical anomaly detection (e.g., MVTec LOCO [171]), pedestrian anomaly detection (e.g., UCF-Crime Dataset [172]), traffic anomaly detection (e.g., Kaggle Accident Detection [173]), and time series anomaly detection (e.g., Outlier Detection Dataset [174]).
Evaluation  Evaluation in anomaly detection with LVLMs is an open challenge. AnomalyGPT [144] asks LVLMs the question “Is there an anomaly in this image?” and determines anomaly or normal based on the simple rule-based approach of whether the response contains a “yes” or “no”. However, this rule-based approach is not robust, as a response is considered correct even if the explanation following “yes” is completely incorrect. On the other hand, Cao et al. [142] conducted only qualitative evaluations and left quantitative evaluations as an open challenge. Therefore, the evaluation of anomaly detection by LVLM is a future challenge.
@@ 6.3.2 Findings
Cao et al. [142] described the observations of GPT-4V in the paper, so we briefly summarized them here.
@@
7 Potential Challenges
This section discusses potential challenges for CLIP-based OOD detection that may be highlighted by the widespread adoption of our framework. Since similar or ambiguous problem settings currently exist, we present these challenges to avoid future confusion for readers.
Figure 5: Illustration of the settings for in-class open-vocabulary OOD detection and zero-shot open-vocabulary OOD detection. In-class open-vocabulary OOD detection should be compared with few-shot OOD detection methods, while zero-shot open-vocabulary OOD detection should be discussed in the context of zero-shot OOD detection.
7.1 Open-vocabulary OOD Detection
The existing setting for open-vocabulary OOD (OV-OOD) detection [47] involves training with images from a small subset of ID classes and evaluating with all ID class names, as explained in Sec. 4.2.1. However, this problem definition has several issues: (i) The OOD detection performance on ID classes not included in the training subset is unclear, (ii) This contradicts the settings of the existing open-vocabulary settings, including CLIP-based AD (Sec. 5.1.2), which assume no explicit class overlap between training and test ID data [175, 176, 63, 29, 65], leading to confusion in the fields. Therefore, for clear evaluation and common understanding in the field, we redefine the OV-OOD detection as in-class open-vocabulary OOD detection (in-class OV-OOD detection) and zero-shot open-vocabulary OOD detection (zero-shot OV-OOD detection). The illustration for each setting is shown in Fig. 5. Detailed explanations for each setting are as follows:
In-class OV-OOD Detection  In-class OV-OOD detection is the same as the original setting proposed by  [47], involving training with images from a small subset of ID classes and evaluating on all ID class names including training classes. In-class OV-OOD detection should be evaluated in the few-shot OOD detection setting to assess how well these methods retain their performance when trained with limited class data. In this setting, it is preferable to employ the existing benchmark setup that randomly selects subset ID classes from all ID classes [47].
Zero-shot OV-OOD Detection  Zero-shot OV-OOD detection shares the training setting with in-class evaluation, using a small subset of ID classes. However, it differs in the testing phase, where it exclusively utilizes classes not included in the training subset. This problem setting is consistent with the common open-vocabulary setting [175, 176] and the open-vocabulary setting in CLIP-AD [63, 29, 65]. Zero-shot OV-OOD detection can be positioned within the area of zero-shot OOD detection, which would facilitate the research of zero-shot OOD detection. When selecting the subset of classes for training, we should ensure semantic dissimilarity between the classes in the evaluation and training sets to guarantee the validity of the zero-shot problem setting. For instance, when splitting ImageNet classes into training and evaluation subsets for zero-shot OV-OOD detection, it would be preferable to consider the hierarchical classes of ImageNet to avoid similar semantics between training and evaluation subsets.
7.2 CLIP-based Full-spectrum OOD Detection
Figure 6: Illustration of the settings for full-spectrum OOD detection, open-set domain generalization, and hard full-spectrum OOD detection. Hard full-spectrum OOD detection is a potential challenge arising from our generalized OOD detection v2.
CLIP-based full-spectrum OOD detection is a well-established task with existing research [115]. However, with the future prevalence of hard OOD detection, which incorporates the OSR setup, the distinction between this task and open-set domain generalization (OSDG)  [177, 178, 179, 180] may become ambiguous. OSDG shares a similar motivation of realizing models capable of both generalization and detection. The illustration of both problem settings is shown in Fig. 6 (a)(b). OSDG assumes that only data with covariate shifts will be input during testing, and aims to classify input samples into one of ID classes if their semantics align with ID classes and detect OOD that exhibit semantic shifts. OSDG is a research area with a much smaller number of works compared to the main topics of this survey (OD, AD, ND, OSR, and OOD detection), but, in recent years, some works have tackled OSDG using CLIP [71, 181, 182].
To eliminate potential ambiguity, this paper defines a hard full-spectrum OOD (FS-OOD) detection, following our generalized OOD detection v2. The illustration of the problem setting for hard FS-OOD detection is shown in Fig. 6 (c). Hard FS-OOD detection expands the scope of existing FS-OOD detection by utilizing the conventional OSR setup. It introduces OOD samples from different categories within the training domain and OOD samples that share the same covariate shift as the target domain but exhibit semantic shifts. This task can be regarded as an extensive version of both existing FS-OOD detection and OSDG tasks. Hard FS-OOD detection represents a developing challenge with promising potential for future research.

### v2

@@ 5.4.2 Similarity between Each Methodology
Each Problem Setting  The existing problem settings for VLM-based AD and VLM-based OOD detection are similar. Both primarily focus on zero-shot and few-shot settings and can be categorized into training-free, auxiliary training-based, and ID training-based methods. Examining each problem more closely provides valuable insights into future directions for both fields.
@@
History of Approaches  The history of the progress of methods for VLM-based AD and VLM-based OOD detection is similar. For instance, both problems initially started with naive methods with manual OOD prompts (ZeroOE (Fort et al., 2021) for OOD detection, WinCLIP (Jeong et al., 2023) for AD). To address the issues with these initial approaches, subsequent methods emerged that replaced OOD prompts with learnable parameters (LSN (Nie et al., 2023) and NegPrompt (Li et al., 2024b) for OOD detection, and AnomalyCLIP (Zhou et al., 2024a) for AD). Therefore, by carefully examining each other’s fields, there is potential for mutual enhancement and interaction in the future
6 Benchmarks and Experiments
In this section, we compare several representative VLM-based OOD detection methods.
6.1 Benchmarks and Evaluation Metrics
In OOD detection, it is common to consider an entire dataset as ID and to use several other datasets that are semantically disjoint from any ID classes as OOD datasets. However, as described in Sec.3.1, recent advances in hard OOD detection focus on constructing benchmarks where both ID and OOD samples are derived from the same large-scale dataset such as ImageNet. This setup is inspired by the benchmark design of OSR.
Following prior work (Zhang et al., 2024b; Ming et al., 2022a), we adopt the following terminology: among OOD samples drawn from datasets different from the ID dataset, those that are semantically similar to the ID classes are referred to as near OOD, while those that are semantically distant are referred to as far OOD. When both ID and OOD samples are derived from the same dataset, as in OSR, we refer to this setting as hard OOD.
In this section, we report results on the widely used ImageNet OOD benchmark (Huang & Li, 2021) and two ImageNet-based hard OOD benchmark.
ImageNet Far OOD Benchmark  In the ImageNet OOD benchmark, ImageNet is used as the ID dataset, while datasets such as iNaturalist (Van Horn et al., 2018) serve as OOD datasets. Among these, the most commonly used benchmark involves using the ImageNet validation set as the ID data and the following four datasets as OOD data: iNaturalist (Van Horn et al., 2018), SUN (Xiao et al., 2010), Places (Zhou et al., 2017), and TEXTURE (Cimpoi et al., 2014). iNaturalist contains approximately 859,000 images of plants and animals across over 5,000 species. For OOD detection evaluation, 10,000 images are randomly sampled from 110 classes that are disjoint from ImageNet-1K. SUN consists of over 130,000 scene images spanning 397 categories. For OOD evaluation, 10,000 images are randomly sampled from 50 categories that do not overlap with ImageNet. Places is another scene-centric dataset with a similar concept space to SUN. For OOD detection, 10,000 images are randomly sampled from 50 non-overlapping classes with ImageNet-1K. TEXTURE comprises 5,640 real-world texture images from 47 categories. The entire dataset is used for OOD evaluation. When training is required, the training set of ImageNet is used for model training.
ImageNet Near OOD Benchmark  In previous studies, two datasets, SSB-Hard (Vaze et al., 2022a) and NINCO (Bitterwolf et al., 2023), have been proposed as OOD datasets that are more semantically similar to ImageNet. SSB-Hard contains 49,000 images from 980 categories, which are selected from ImageNet-21K (Ridnik et al., 2021). NINCO is a dataset with 5,879 manually collected images (Bitterwolf et al., 2023), which is constructed to be semantically close to ImageNet-1K, but without any overlapping classes.
ImageNet-20 OOD Benchmark  The ImageNet-20 OOD Benchmark is one of the earliest benchmarks proposed for hard OOD detection (Ming et al., 2022a). In this setting, ImageNet-20 is used as the ID dataset, and ImageNet-10, which has no overlapping categories, is used as the OOD dataset. The direction can also be reversed. These subsets are selected with reference to CIFAR-10 (Krizhevsky et al., 2009a) classes, making the semantic distance between ImageNet-20 and ImageNet-10 relatively small. ImageNet-20 contains 1,000 images, while ImageNet-10 contains 500 images.
ImageNet-X OOD Benchmark  ImageNet-X is a recently proposed benchmark for hard OOD detection (Noda et al., 2025). It leverages the semantic hierarchy of ImageNet and creates ID and OOD splits within the same superclass. Both ID and OOD sets consist of 500 classes each. The ID and OOD subsets of ImageNet-X contain 25,000 images respectively.
6.2 Metrics
We use the ID accuracy and AUROC scores. AUROC scores measure the area under the Receiver Operating Characteristic (ROC) curve.
6.3 Experimental Setup
We compare commonly used eight OOD detection methods in zero-shot and few-shot settings.
Zero-shot Methods  For the zero-shot setting, we use five methods: MCM (Ming et al., 2022a), GL-MCM (Miyai et al., 2025b), NegLabel (Jiang et al., 2024), EOE (Cao et al., 2024a), and CLIPN (Wang et al., 2023a). As shown in Table 2, MCM and GL-MCM are categorized as methods without training and without OOD prompts. NegLabel and EOE are categorized as methods without training but with OOD prompts. CLIPN is categorized as a method with training and with OOD prompts.
Few-shot Methods  For the few-shot settings, we use CoOp (Zhou et al., 2022c; Ming & Li, 2024a), LoCoOp (Miyai et al., 2023b), and ID-like-Prompt (IDPrompt) (Bai et al., 2024a). CoOp and LoCoOp use MCM as the detection method during inference. For CoOp and LoCoOp, we follow the hyperparameter settings from previous studies and train with 16 shots. On the other hand, since IDPrompt requires higher training costs, we follow the original implementation (Bai et al., 2024a) and conduct training with only 1 shot.
Table 3: Comparison of OOD detection methods across ImageNet, ImageNet-20, and ImageNet-X. We use AUROC for the evaluation of OOD detection.
Method Train Type OOD Prompts ImageNet ImageNet-20 ImageNet-X
Far-OOD Near-OOD ID Acc. Hard-OOD ID Acc. Hard-OOD ID Acc.
Zero-shot
MCM Free 90.66 68.91 66.73 97.38 91.30 74.30 75.97
GL-MCM Free 91.47 71.21 66.73 98.49 91.30 75.53 75.97
NegLabel Free 94.15 73.33 66.73 97.47 91.30 72.15 75.97
EOE Free 92.95 70.22 66.73 98.16 91.30 77.07 75.97
CLIPN-A Auxiliary 94.11 80.94 66.73 97.16 91.30 77.08 75.97
Few-shot
CoOp ID 91.16 76.46 71.77 97.78 94.80 75.23 80.38
LoCoOp ID 93.02 69.11 71.61 99.04 94.67 75.17 80.42
ID-like-Prompt ID 92.46 67.18 68.43 92.90 89.80 66.01 77.18
6.4 Experimental Results
The experimental results are summarized in Table 3. Below, we highlight several key findings.
Performance Rankings Vary Across Datasets  From Table 3, we observe that the ranking of methods differs significantly depending on the dataset. For example, LoCoOp outperforms CoOp on the standard ImageNet far OOD benchmark, but underperforms on the near OOD benchmarks. This suggests that LoCoOp’s OOD regularization strategy, which uses background OOD images during training, is especially effective for detecting semantically distant classes such as in far OOD settings, but not effective for near OOD settings. These results indicate the importance of evaluating OOD detection methods across multiple benchmarks, rather than relying on a single one.
ImageNet-20 May Already Be Saturated  On ImageNet-20, most methods achieve over 97% accuracy, making it difficult to distinguish performance differences among methods. Therefore, future research on OOD detection is likely to shift toward benchmarks with more diverse and larger numbers of classes.
Room for Improvement Remains in Near and Hard-OOD Detection  The results on ImageNet near OOD and ImageNet-X show that the methods with the highest score reach only around 80% accuracy. This indicates significant room for further improvement. We believe that future research in OOD detection will increasingly focus on these challenging benchmarks.
@@ 7 Evolution in LVLM Era
In this section, we introduce the early advances in OOD detection and AD in the Large Vision Language Models (LVLM) era. While previous sections focused on VLMs such as CLIP, this section shifts our focus to the more emerging topic of “Large” VLMs. Recent advancements in computer vision have led to the emergence of LVLMs such as GPT-4V (OpenAI et al., 2023) and LLaVA (Liu et al., 2024b). Although these fields are still in their early stages with limited papers, this survey provides a deep introduction to each problem in the hope that our detailed review can help foster further advancements in this area.
Figure 6: Overview of the evolution of each problem in the Large Vision Language Model (LVLM) era. (i) Sensory AD has consistently been an active research field in the LVLM era (Cao et al., 2023b). (ii) Notably, OOD detection is evolving into a new task called Unsolvable Problem Detection (Miyai et al., 2025a) in the LVLM era. Figure adapted partially from (Cao et al., 2023b) and from (Miyai et al., 2025a).
@@ 7.1 Change of Each Problem
i. Sensory AD $\rightarrow$ Sensory AD  Sensory AD has continued to develop in the LVLM era (Gu et al., 2024b; Cao et al., 2023b; Li et al., 2023b). The use of LVLMs has made AD applicable in many domains and modalities (Cao et al., 2023b).
ii. OOD Detection $\rightarrow$ Unsolvable Problem Detection  In the LVLM era, OOD detection has evolved into a new task termed Unsolvable Problem Detection (UPD) (Miyai et al., 2025a). UPD evaluates the LVLMs’ ability to recognize and abstain from answering unexpected or unsolvable input questions, effectively expanding the scope of OOD detection into the context of Visual Question Answering (VQA) tasks. This shift to the VQA task has significantly broadened the concept of OOD detection to a wider range of AI tasks involving LVLMs.
@@ 7.2.1 Summary of Problem
Background  Following the recent revolutionary development of LLMs (Chiang et al., 2023; Touvron et al., 2023; Wei et al., 2023; Zhao et al., 2023), LVLMs (Awadalla et al., 2023; Dai et al., 2023; Liu et al., 2024b; Wang et al., 2023b; Ye et al., 2023; Li et al., 2023a; Lin et al., 2024) have demonstrated remarkable capabilities in diverse applications (Liu et al., 2024a; f; Yue et al., 2024). However, a significant concern has arisen regarding the reliability of these models, specifically their ability to generate accurate and trustworthy information. These models frequently produce incorrect or misleading information, a phenomenon referred to as “hallucination” (Bai et al., 2024b). Among the various hallucination issues (Bai et al., 2024b), the challenge of identifying out-of-place questions is crucial for deploying LVLMs in safety-critical applications. This challenge extends the concept of OOD detection to the VQA tasks for LVLMs and represents a specific aspect of LVLMs’ trustworthiness.
@@
Definition  Unsolvable Problem Detection (UPD) is a task to measure the trustworthiness of LVLMs, which is designed to evaluate models’ capacity to withhold answers when faced with unsolvable problems. The UPD task can be categorized into three distinct problem types: Absent Answer Detection (AAD), Incompatible Answer Set Detection (IASD), and Incompatible Visual Question Detection (IVQD). Examples in each setting are shown in the right of Fig. 6. AAD evaluates the model’s capacity to determine when the correct answer is absent from the provided options. IASD assesses the model’s ability to discern answer choices that are completely irrelevant to the given question and image. IVQD evaluates the model’s capacity to discern whether a question and image are unrelated or mismatched.
@@
BenchmarkMiyai et al. (2025a) created MM-UPD Bench for the UPD challenge. MM-UPD encompasses MM-AAD, MM-IASD, and MM-IVQD benchmarks for each UPD problem. Each benchmarks are created on the top of MMBench (dev) (Liu et al., 2024f), which is a systematically designed objective benchmark for evaluating various abilities of LVLMs. Following the definition of each ability in MMBench (e.g., “Coarse Perception: Image Scene” and “Logic Reasoning: Future Prediction”), MM-UPD evaluates the trustworthiness of LVLMs from various abilities.
@@
Although MM-UPD is the main benchmark, the adaptation cost of creating UPD problems is not high, making it highly applicable to other benchmarks. For instance, the recently proposed MuirBench (Wang et al., 2025), a benchmark for multi-image understanding, has incorporated the concept of UPD by adding unsolvable problems.
@@
Application  UPD has a wide range of applications, from everyday use of LVLMs to robot manipulation. Especially when incorporating LVLMs into safety-critical domains such as robot manipulation (Liu et al., 2024d) and autonomous driving (Li et al., 2024d), there is a risk of significant problems if the LVLM fails to identify erroneous user questions and makes incorrect predictions. UPD serves as a task to ensure safety in such safety-critical scenarios.
Evaluation  UPD employs three evaluation metrics: Standard Accuracy, which measures performance on solvable problems; UPD Accuracy, which assesses accuracy on unsolvable problems; and Dual Accuracy, which considers both aspects. Dual Accuracy considers a response correct only if the model successfully answers both the solvable and unsolvable problems within each paired set. The rationale is that ideal LVLMs should not only give correct answers for the solvable problems but need to withhold answering for unsolvable problems.
@@ 7.2.2 Findings
In the following, we briefly summarized the findings of the UPD challenge (Miyai et al., 2025a).
1. Most LVLMs Hardly Hesitate to Answer. Most LVLMs, especially open-source LVLMs, have significantly low UPD accuracies, which indicates the difficulty of the UPD challenge. For example, LLaVA-1.5 (Liu et al., 2024b) and CogVLM (Wang et al., 2023b), which are state-of-the-art LVLMs, completely fail to withhold answering. GPT-4V achieves higher performances than other LVLMs due to its safety training process (OpenAI, 2023). However, there is still a performance gap from the upper bound scores.
@@ 7.3.1 Summary of Problem
Background  Anomaly detection is a crucial task in a variety of domains and data types. However, existing anomaly detection models are often designed for specific domains or modalities (Cao et al., 2023b). Also, current AD methods only provide an anomaly score for the test sample and require a manual threshold to distinguish between normal and anomalous instances for each sample (Gu et al., 2024b). To facilitate real-world applications, developing a system capable of expressing anomalies in natural language across various modalities and domains is crucial for ensuring accessibility to a wider range of users.
Definition  The definition of AD remains consistent with conventional and VLM-based AD, aiming to identify samples that deviate from predefined normality. The key difference lies in the output format. Traditional methods generate an anomaly score as a numerical value, requiring a manual threshold to determine whether a sample is anomalous. In contrast, AD with LVLMs aims to recognize and describe anomalies using text, removing the need for manual thresholds and improving human interpretability.
Benchmark  Since the field of AD with LVLMs is in its infant stage, there are still no unified benchmarks. AnomalyGPT (Gu et al., 2024b) focuses on industrial image anomaly detection/localization and uses the standard benchmarks MVTec-AD (Bergmann et al., 2019) and VisA (Zou et al., 2022). More recently, Cao et al. (2023b) extend the domain and modality and demonstrate the applications in industrial image anomaly detection/localization (e.g., MVTec-AD (Bergmann et al., 2019)), point cloud anomaly detection (MVTec 3D (Bergmann et al., 2022b)), medical image anomaly detection/localization (e.g., Chest X-ray (Kermany et al., 2018), Head CT (Felipe, 2018)), logical anomaly detection (e.g., MVTec LOCO (Bergmann et al., 2022a)), pedestrian anomaly detection (e.g., UCF-Crime Dataset (Sultani et al., 2018)), traffic anomaly detection (e.g., Kaggle Accident Detection (Kay, 2018)), and time series anomaly detection (e.g., Outlier Detection Dataset (User, 2021)).
Evaluation  Evaluation in anomaly detection with LVLMs is an open challenge. AnomalyGPT (Gu et al., 2024b) asks LVLMs the question “Is there an anomaly in this image?” and determines anomaly or normal based on the simple rule-based approach of whether the response contains a “yes” or “no”. However, this rule-based approach is not robust, as a response is considered correct even if the explanation following “yes” is completely incorrect. On the other hand, Cao et al. (2023b) conducted only qualitative evaluations and left quantitative evaluations as an open challenge. Therefore, the evaluation of anomaly detection by LVLM is a future challenge.
@@ 7.3.2 Findings
Cao et al. (2023b) described the observations of GPT-4V in the paper, so we briefly summarized them here.